Shift Summary, Day shift 11 Sep 1999

Tony Barker (tonyb@cuhep.Colorado.EDU)
Sat, 11 Sep 1999 14:40:59 -0600 (MDT)

Shift Summary for Day Shift, 09/11/99

On Shift: Bob Hsiung, Tony Barker

Hours of beam available at all intensities: 8.50
Hours of beam available at nominal intensity: 8.50
Hours of E832 data (special runs not included): 8.00
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time: Replace Etot board,
reset fruits, change tapes.

========================================================================

Runs ended this shift: 14500 7.25M events; banana readout problems.

Tapes brought to FCC this shift: QKK411-413
QKS411-413
QKC411-413
QKB145

========================================================================

Accesses during shift: At start of shift, to replace Etot board.

========================================================================

Overview of shift: 832 data taking at 4.0 E12

Sydney and Albert replace Etot board to fix
inefficient bit for CsI channel 920.

Blockhouse dehumidifier is broken.
Humidity creeping up past 10%; Ron Ray
is paged, calls Therm Flow, which is
responsible for the servicing of the
dehumidifier. Their number is 847-541-0029,
and the usual service tech is Larry Grimshaw.
Larry arrives about 14:30 and promptly fixes
the BryAir.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14500 001 07:10 Start run prior to access to replace Etot board.

07:30 Access finished; open beam stop.

177 11:08 See trigger problems in banana and YTF plots;
decide to stop run and start a new one.

14501 001 11:20 Update CsI and FERA constants as 14501 starts,
also reboot FSCC's and reset fruits.
Previous banana problems seem to have been
in the readout (so that the run 14500 data
is otherwise ok). Problems are gone in 14501.

094 13:25 Blockhouse humidity alarms have been going
of for awhile. Ron Ray paged; is here trying
to contasct someone (anyone!) who can come in
to repair the dehumidifier. It appears that the
dehumidifier's desiccant heater is not working.
BlckH-Up humidity is currently 9.8%; BlckH-Dn
is 12.0%.

122 14:05 Error messages from online split:
'QKA027: Cache control file does not exist'

157 14:50 Larry Grimshaw finishes repairs to BryAir
dehumidifier. Humidity drops rapidly.

168 15:05 Alarm: CsI 2534 has 10% laser rms shift and
bad capid alarms. Several other channels
show shifts of 3-5%. Possibly due to high
electronics temperature resulting from use
of backup dehumidifier.

187 15:30 Shift ends.

========================================================================
Follow up Notes on Previous Shift

CsI and FERA constants were updated at beginning of run 14501.

========================================================================
Items to Follow up on Next Shift

Monitor HCC hotbits and inefficient channels carefully.
High temperature and multiple power cycling during
9/10 day shift may have caused some damage.

Monitor laser stability rms for bad channels, especially 2534.

If humidity alarms go off (humidity above 10% in the blockhouse)
call Leonard Nelson. His home number is 630-231-1416; pager
888-750-0702.